# Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning

### v1

@@ Abstract
Teaching large language models (LLMs) to be faithful in the provided context is crucial for building reliable information-seeking systems. Therefore, we propose a systematic framework, Canoe, to improve the faithfulness of LLMs in both short-form and long-form generation tasks without human annotations. Specifically, we first synthesize short-form question-answering (QA) data with four diverse tasks to construct high-quality and easily verifiable training data without human annotation. Also, we propose Dual-GRPO, a rule-based reinforcement learning method that includes three tailored rule-based rewards derived from synthesized short-form QA data, while simultaneously optimizing both short-form and long-form response generation. Notably, Dual-GRPO eliminates the need to manually label preference data to train reward models and avoids over-optimizing short-form generation when relying only on the synthesized short-form QA data. Experimental results show that Canoe greatly improves the faithfulness of LLMs across 11 different downstream tasks, even outperforming the most advanced LLMs, e.g., GPT-4o and OpenAI o1.111 The data, code, and models will be available at https://github.com/S1s-Z/CANOE.
[…]
tcb@breakable
Teaching Large Language Models to Maintain Contextual Faithfulness via Synthetic Tasks and Reinforcement Learning
Shuzheng Sithanks:  Equal Contribution.♠♢, Haozhe Zhao11footnotemark: 1, Cheng Gao11footnotemark: 1, Yuzhuo Bai, Zhitong Wang Bofei Gao, Kangyang Luo, Wenhao Li, Yufei Huang, Gang Chen Fanchao Qi, Minjia Zhang, Baobao Chang, and Maosong Sun Tsinghua University   Peking University   DeepLang AI University of Illinois Urbana-Champaign
@@ 1 Introduction
Recent progress in large language models (LLMs) has revolutionized text generation with their remarkable capabilities (OpenAI, 2023; DeepSeek-AI et al., 2025b). LLMs are widely used to generate fluent and coherent text responses based on the provided contextual information, e.g., document question answering (QA) (Wang et al., 2024) and text summarization (Zhang et al., 2024). However, LLMs often generate responses that are not faithful or grounded in the input context, i.e., faithfulness hallucinations (Ji et al., 2023; Huang et al., 2024; Si et al., 2025), which can undermine their trustworthiness. Maintaining faithfulness to the context is especially important in fields where accurate information transfer is essential (Duong et al., 2025). For instance, in legal summarization (Dong et al., 2025), the text output must reflect the content of legal documents without introducing any distortions.
@@
Figure 1: Average score on 11 downstream tasks vs model size. With only 7B parameters, Canoe already exceeds state-of-the-art LLMs like GPT-4o and o1.
@@
However, improving the faithfulness of LLMs faces three key challenges. Specifically, (1) Faithfulness is difficult to improve by simply scaling model parameters: Previous works (Xie et al., 2024; Li et al., 2025) find that LLMs may overly rely on internal knowledge learned from extensive pre-training data while disregarding provided contexts, i.e., the knowledge conflicts (Xu et al., 2024b). When the model parameters increase and internal knowledge grows, this may lead to greater knowledge conflicts and further lower the faithfulness of LLMs (Ming et al., 2025). Thus, it is necessary to explore the tailored post-training method to improve the faithfulness instead of simply scaling the model parameters. (2) Faithfulness is challenging to consistently boost across different downstream tasks: Recently, several methods (Li et al., 2024; Duong et al., 2025) have been proposed to improve the faithfulness of LLMs for different tasks. For example, Bi et al. (2024) aligns LLMs through DPO (Rafailov et al., 2023) with constructed faithful and unfaithful short-form completions, improving the performance of LLMs on short-form QA tasks. However, these recent methods are designed for specific tasks, so they fail to consistently improve the faithfulness of LLMs across various tasks, like text summarization and multiple-choice questions, because these tasks can vary greatly. (3) Data used to enhance faithfulness is hard to scale: This issue is especially problematic with data used to improve the faithfulness in long-form generation tasks. Unlike tasks with clear answers, e.g., short-form fact-seeking QA tasks (Wei et al., 2024), there is no standard way to ensure data quality in long-form generation tasks (Duong et al., 2025). Thus, data is typically annotated by humans (Kryscinski et al., 2020; Zhu et al., 2023), which is costly and not scalable.
@@
To tackle these challenges, we propose a systematic post-training method called Canoe. The main idea behind Canoe is to synthesize easily verifiable short-form QA data and then leverage reinforcement learning (RL) with tailored rule-based rewards to improve the faithfulness of LLMs in both short-form and long-form generation tasks. Canoe firstly introduces Dual-GRPO, a variant of GRPO (Shao et al., 2024) that includes three carefully tailored rule-based RL rewards derived from synthesized short-form QA data, while optimizing both short-form and long-form response generation. For the provided contextual information and question, Dual-GRPO first prompts LLMs to produce a reasoning process, followed by a long-form answer composed of detailed and complete sentences, and finally a concise short-form answer in just a few words. In this way, we can assign different rewards to long-form and short-form responses, optimizing both simultaneously. Note that we assign accuracy rewards on generated short-form responses since the short-form QA task enables reliable rule-based verification of faithfulness. To overcome the problem of the faithfulness of the generated long-form responses being difficult to evaluate via rule-based verification (Zheng et al., 2025; OpenAI, 2025), we propose proxy rewards to evaluate it implicitly. Specifically, we construct the new input by replacing the given context with the generated long-form answer, then feed it to the LLMs to evaluate whether a long-form answer can drive the LLMs toward the correct short-form answer. If the generated long-form response enables LLMs to generate the correct final answer, this indicates that it remains context-faithful and contains easy-to-understand sentences that answer the question correctly. We also introduce format rewards to ensure more structured outputs and contribute to more stable training. To obtain the data used for training without human annotation, we collect head-relation-tail triples from the knowledge base, apply the advanced GPT-4o (OpenAI, 2023) to synthesize the question and contextual information, and use the tail entity from the triple as the answer to ensure the correctness. Moreover, we introduce four diverse QA tasks to ensure the complexity and diversity of the training data. Combined with the rule-based Dual-GRPO and data synthesis, Canoe can teach LLMs to remain context-faithful in both short-form and long-form generation tasks without relying on human annotations.
@@
We evaluate the effectiveness of Canoe across 11 different downstream tasks, covering short-form and long-form generation tasks. Results show that Canoe significantly reduces faithfulness hallucinations. Specifically, Canoe significantly improves the overall score, e.g., 22.6% for Llama3-Instruct-8B. Meanwhile, Canoe surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score. To the best of our knowledge, these results are unprecedented for open-source models that do not rely on additional human annotations.
2 Related Work
Recently, the demand for utilizing LLMs to generate coherent text responses based on the provided contexts has continued to grow, particularly in text summarization and retrieval-augmented generation (RAG) scenarios. However, LLMs are often criticized for generating outputs that deviate from the provided contents, namely faithfulness hallucination (Li et al., 2022; Ji et al., 2023; Si et al., 2023; Huang et al., 2024). Many approaches have been proposed to improve the faithfulness of LLMs. The first line of work focuses on the inference stage of LLMs, such as designing prompts to encourage context integration (Zhou et al., 2023), improving context quality via explicit denoising (Xu et al., 2024a), and context-aware decoding to amplify contextual information (Shi et al., 2024). Although effective, these approaches primarily serve as a compensatory way rather than enabling the model to inherently learn to prevent generating unfaithful responses. Therefore, many studies attempt to apply post-training methods to improve the faithfulness. Bi et al. (2024) utilizes constructed faithful and unfaithful short-form completions and applies DPO to align LLMs to be context-faithful in short-form QA tasks. Huang et al. (2025) trains LLMs to discriminate between faithful and unfaithful responses in long-form QA tasks by unfaithful response synthesis and contrastive tuning. Duong et al. (2025) proposes a pipeline to generate a self-supervised task-specific dataset and applies preference training to enhance the faithfulness for a special task. However, these methods struggle to consistently improve the faithfulness of LLMs across various tasks, as these methods are designed for specific tasks. Thus, how to consistently improve the faithfulness of LLMs on different downstream tasks, including short-form and long-form generation tasks, still remains under-explored.
@@ 3.1 Training Data Construction
Constructing high-quality and easily verifiable data is crucial for rule-based RL training (Shao et al., 2024). Inspired by knowledge base question generation (Cui et al., 2019; Guo et al., 2024), we attempt to collect triples from the knowledge base and use the advanced LLMs to synthesize the context and question. Concretely, we first collect about 30,000 head-relation-tail triples from Wikidata (Vrandečić and Krötzsch, 2014). Each collected triple $(h,r,t)$ includes a head entity $h$, a tail entity $t$, and the relation $r$ between two entities. Then we craft prompt templates and query the most advanced GPT-4o to synthesize the contextual information $c$ and question $q$ based on the triple $(h,r,t)$. We directly use the tail entity $t$ as the final answer $a$ to ensure the correctness and easy validation of the synthesized data. Each synthetic short-form QA sample $(c,q,a)$ consists of a contextual passage $c$, a question $q$, and a ground truth answer $a$. In this way, we can obtain short-form QA data that can be easily verified, thus we can utilize a rule-based RL method to optimize our LLMs to be more faithful. Meanwhile, to ensure the complexity and diversity of training data, we design four diverse QA tasks, including straightforward context, reasoning-required context, inconsistent context, and counterfactual context. The model is expected to answer the question by leveraging the information in the provided context.
@@
Reasoning-required Context. This context contains multiple related entities and relations, and requires models to answer multi-hop reasoning questions. Firstly, we construct a subgraph based on the sampled triples and extract 2, 3, 4-hop paths $[(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and employ the constructed paths to query GPT-4o to obtain the multi-hop context and question.
@@
By introducing four different tasks, we construct 10,000 QA pairs used for training without human annotation. These short-form QA data can be easily verified and include tasks varying in complexity, which can make rule-based RL training more efficient in improving the faithfulness of LLMs. More details can be found in the Appendix A, e.g., used prompts, data mixing recipes, and data statistics.
@@ 3.2 Reinforcement Learning Protocol
For RL training of LLMs, methods based on policy optimization, such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024), have been explored. Given the effectiveness of GRPO in training models and its advantages over PPO, e.g., eliminating the need for human-annotated preference data to train a reward model, we utilize GRPO to optimize and improve the faithfulness of the policy model $\pi_{\theta}$.
@@
$\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{c,q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref})\right],\end{aligned}$
$\mathcal{L}_{i}&=\min\left(w_{i}A_{i},\text{clip}(w_{i},1-\epsilon,1+\epsilon)A_{i}\right),\end{aligned}$
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$, $\pi_{\theta_{\text{old}}}$ is the policy before the update, $\pi_{\text{ref}}$ is the reference policy (i.e., the initial model), $\epsilon$ and $\beta$ are hyperparameters controlling the update step and divergence regularization and $A_{i}$ is computed using the normalized reward within the group. We use synthesized short-form QA data as training data, which is easily verifiable, so that we can apply GRPO and train LLMs using the rule-based reward function. By generating multiple candidates per input, GRPO naturally accommodates the inherent challenges of utilizing the contextual information $c$ and answering the question $q$, e.g., LLMs may overly rely on the internal knowledge while disregarding provided contexts. Meanwhile, employing the rule-based GRPO removes the need for humans to annotate short-form and long-form preference data used for training the reward model.
@@ 3.3 Reward Design
Having a well-designed reward is key to the effectiveness of RL training (Du et al., 2025). To use easily verifiable short-form QA data to improve the faithfulness, the most intuitive reward would be the accuracy reward, which can check if the generated responses match the ground truth answers. However, in our early experiments, we found that relying solely on short-form QA data and accuracy rewards fails to enhance the faithfulness of long-form response generation, as the models may over-optimize short-form generation and learn a false pattern. For example, the tuned models tend to simply copy text spans from the context as answers and lose their ability to generate long-form responses. Unfortunately, directly evaluating the faithfulness of long, free-form responses via the rule-based verification continues to pose a significant and unresolved challenge.
Therefore, we propose Dual-GRPO, which includes a set of well-designed rewards that provide more harmonized guidance for optimizing LLMs to generate faithful responses. Unlike the original GRPO that over-optimizes short-form generation, we first prompt LLMs to generate both long-form and short-form responses, then assign different rewards to the two generated responses to improve the faithfulness of the two types of generation.
System Prompt and Rollouts. For the provided context and question, Dual-GRPO employs the designed system prompt that requires LLMs to produce a reasoning process, then a long-form answer composed of detailed and complete sentences, and finally a concise short-form answer in just a few words. For example, given the context, if the question is “What is the country of origin of Super Mario?”, the long answer could be “Super Mario originated from Japan.”, while the short answer could simply be “Japan”. In this way, we can assign different reward scores to long-form and short-form answers while optimizing them both at once. This system prompt also triggers zero-shot chain-of-thought reasoning in the policy model, which progressively improves as training advances to optimize for the reward. The system prompt used for Dual-GRPO rollouts is shown in the Appendix B.
@@
Proxy Reward for Long-form Response Generation. Evaluating the faithfulness of the generated long-form responses via the rule-based verification remains challenging. This is because these long-form answers are often free-form, making rule-based verification ineffective (Zheng et al., 2025; OpenAI, 2025). Therefore, instead of directly evaluating the faithfulness of the long-form response, we propose a proxy reward to evaluate it implicitly, as the faithfulness of a long-form answer can be measured by its ability to drive the LLMs toward a correct short-form answer. Specifically, for each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and feed it to the LLM to check whether the LLM can produce the correct short-form answer based on $y_{lf}$. If the generated long-form response can enable the LLM to generate the correct answer, it indicates that the long-form response stays faithful to the context, contains complete and easy-to-understand sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct final answer, and a reward score of 0 for those that lead to an incorrect answer.
Format Reward. We also include a format reward that encourages adherence to a predefined output structure (e.g., using <think>, <long_answer>, and <short_answer> tags). Outputs that conform to this pattern receive a reward boost, thereby enhancing clarity and consistency. We use the string matching method to evaluate whether the generated responses adhere to the format, giving a score of 1 for a match and 0 for a mismatch.
Finally, we use the sum of these three rewards as the final composite reward. It enhances the efficacy of the rule-based RL training framework, guiding the model toward generating more faithful responses in both short-form and long-form tasks. More details are shown in the Appendix B.

### v3

@@ Abstract
Teaching large language models (LLMs) to be faithful in the provided context is crucial for building reliable information-seeking systems. Therefore, we propose a systematic framework, Canoe, to reduce faithfulness hallucinations of LLMs across different downstream tasks without human annotations. Specifically, we first synthesize short-form question-answering (QA) data with four diverse tasks to construct high-quality and easily verifiable training data without human annotation. Also, we propose Dual-GRPO, a rule-based reinforcement learning method that includes three tailored rule-based rewards derived from synthesized short-form QA data, while simultaneously optimizing both short-form and long-form response generation. Notably, Dual-GRPO eliminates the need to manually label preference data to train reward models and avoids over-optimizing short-form generation when relying only on the synthesized short-form QA data. Experimental results show that Canoe greatly improves the faithfulness of LLMs across 11 different tasks, even outperforming the most advanced LLMs, e.g., GPT-4o and OpenAI o1. 111 The data, code, and models will be available at https://github.com/S1s-Z/CANOE. Email: ssz24@mails.tsinghua.edu.cn.
@@ 1 Introduction
Recent progress in large language models (LLMs) has revolutionized text generation with their remarkable capabilities (OpenAI2023GPT4TR). In practice, LLMs are widely used to generate fluent and coherent text responses based on the provided contextual information, e.g., document question answering (QA) (wang-etal-2024-leave) and text summarization (zhang2024systematicsurveytextsummarization). However, LLMs often generate responses that are not faithful or grounded in the input context, i.e., faithfulness hallucinations (hit-survey; si-etal-2025-aligning), which can undermine their trustworthiness. Maintaining faithfulness to the context is especially important in fields where accurate information transfer is essential (duong2025scope). For instance, in legal summarization (dong-etal-2025-termdiffusum), the text output must reflect the content of legal documents without introducing any distortions.
@@
Figure 1: Average score on 11 downstream tasks vs model size. With only 7B parameters, Canoe already exceeds state-of-the-art LLMs like GPT-4o and OpenAI o1.
@@
However, improving the faithfulness of LLMs faces three key challenges. Specifically, (1) Faithfulness is difficult to improve by simply scaling model parameters: Previous works (xie2024adaptive; li2025generatediscriminateevolveenhancing) find that LLMs may overly rely on internal knowledge learned from extensive pre-training data while disregarding provided contexts, i.e., the knowledge conflicts (xu-etal-2024-knowledge-conflicts). When the model parameters increase and internal knowledge grows, this may lead to greater knowledge conflicts and further lower the faithfulness of LLMs (ming2025faitheval). Thus, it is necessary to explore a tailored post-training method to improve the faithfulness instead of simply scaling the model parameters. (2) Faithfulness is challenging to consistently boost across different downstream tasks: Recently, several methods (li-etal-2024-improving-faithfulness; duong2025scope) have been proposed to improve the faithfulness of LLMs for different tasks. For example, bi2024contextdpoaligninglanguagemodels aligns LLMs through DPO (dpo) with constructed faithful and unfaithful short-form completions, improving the performance of LLMs on short-form QA tasks. However, these recent methods are designed for specific tasks, so they fail to consistently improve the faithfulness of LLMs across various tasks, like text summarization and multiple-choice questions, because these tasks can vary greatly.
(3) Data used to enhance faithfulness is hard to scale: This issue is especially problematic with data used to improve the faithfulness in long-form generation tasks. Unlike tasks with clear answers, e.g., short-form fact-seeking QA tasks (wei2024measuringshortformfactualitylarge), there is no standard way to ensure data quality in long-form generation tasks (duong2025scope). Thus, data is typically annotated by humans (zhu-etal-2023-annotating), which is costly and not scalable.
@@
To tackle these challenges, we propose a systematic post-training method called Canoe. The main idea behind Canoe is to synthesize easily verifiable short-form QA data and then leverage reinforcement learning (RL) with tailored rule-based rewards to improve the faithfulness of LLMs in both short-form and long-form generation tasks. Canoe firstly introduces Dual-GRPO, a variant of GRPO (shao2024deepseekmathpushinglimitsmathematical) that includes three carefully tailored rule-based RL rewards derived from synthesized short-form QA data, while optimizing both short-form and long-form response generation. For the provided contextual information and question, Dual-GRPO first prompts LLMs to produce a reasoning process, followed by a long-form answer composed of detailed and complete sentences, and finally a concise short-form and easily verifiable answer in just a few words. For example, given the context, if the question is “What is the country of origin of Super Mario?”, the long answer could be “Super Mario originated from Japan.”, while the short answer could simply be “Japan”. In this way, we can assign different rewards to long-form and short-form responses, optimizing both simultaneously. Note that we assign accuracy rewards on generated short-form responses since the short-form QA task enables reliable rule-based verification of faithfulness. To overcome the problem of the faithfulness of the generated long-form responses being difficult to evaluate via rule-based verification (openai2025deep), we propose proxy rewards to evaluate it implicitly. Specifically, we construct the new input by replacing the given context with the generated long-form answer, then feed it to the LLMs to evaluate whether a long-form answer can drive the LLMs toward the correct short-form answer. If the generated long-form response enables LLMs to generate the correct final answer, this indicates that it remains context-faithful and contains easy-to-understand sentences that answer the question correctly. We also introduce format rewards to ensure more structured outputs and contribute to more stable training. To obtain the data used for training without human annotation, we collect head-relation-tail triples from the knowledge base, apply the advanced GPT-4o (OpenAI2023GPT4TR) to synthesize the question and contextual information, and use the tail entity from the triple as the answer to ensure the correctness. Moreover, we introduce four diverse QA tasks to ensure the complexity and diversity of the training data. Combined with the rule-based Dual-GRPO and data synthesis, Canoe can teach LLMs to remain context-faithful in both short-form and long-form generation tasks without relying on human annotations.
@@
We evaluate the effectiveness of Canoe across 11 different downstream tasks, covering short-form and long-form generation tasks. Results show that Canoe significantly reduces faithfulness hallucinations. Specifically, Canoe significantly improves the overall score, e.g., 22.6% for Llama3-Instruct-8B. Meanwhile, Canoe surpasses the most advanced LLMs (e.g., GPT-4o) in the overall score. As shown in Figure 1, these results are unprecedented for open-source models that do not rely on additional human annotations.
@@ 2.1 Training Data Construction
Constructing high-quality and easily verifiable data is crucial for rule-based RL training (shao2024deepseekmathpushinglimitsmathematical). Inspired by knowledge base question generation (cui-etal-2019-kb; guo-etal-2024-diversifying), we attempt to collect triples from the knowledge base and use the advanced LLMs to synthesize the context and question. Concretely, we first collect about 30,000 head-relation-tail triples from Wikidata (wikidata). Each collected triple $(h,r,t)$ includes a head entity $h$, a tail entity $t$, and the relation $r$ between two entities. Then we craft prompt templates and query the most advanced GPT-4o to synthesize the contextual information $c$ and question $q$ based on the triple $(h,r,t)$. We directly use the tail entity $t$ as the final answer $a$ to ensure the correctness and easy validation of the synthesized data. Each synthetic short-form QA sample $(c,q,a)$ consists of a contextual passage $c$, a question $q$, and a ground truth answer $a$. In this way, we can obtain short-form QA data that can be easily verified; thus, we can utilize a rule-based RL method to optimize our LLMs to be more faithful. Meanwhile, to ensure the complexity and diversity of training data, we design four diverse QA tasks, including straightforward context, reasoning-required context, inconsistent context, and counterfactual context. The model is expected to answer the question by leveraging the information in the provided context.
@@
Reasoning-required Context. This type of context contains multiple related entities and relations, and requires models to correctly answer multi-hop reasoning questions. Firstly, we construct a subgraph based on the sampled triples and extract 2, 3, 4-hop paths $(h^{1},r^{1},t^{1}),...,(h^{n},r^{n},t^{n})]_{n\leq 4}$. Then, we use the $n$-th tail entity $t^{n}$ as the ground truth answer and employ the constructed paths to query GPT-4o to obtain the multi-hop context and question.
@@
By introducing four different tasks, we construct 10,000 QA pairs used for training without human annotation. These short-form QA data can be easily verified and include tasks varying in complexity, which can make rule-based RL training more efficient in improving the faithfulness. More details are shown in the Appendix A, e.g., used prompts, data mixing recipes, and data statistics.
@@ 2.2 Reinforcement Learning Protocol
For RL training of LLMs, methods based on policy optimization, such as PPO (ppo) and GRPO (shao2024deepseekmathpushinglimitsmathematical), have been explored. Given the effectiveness of GRPO in training models and its advantages over PPO, e.g., eliminating the need for human-annotated preference data to train a reward model, we utilize GRPO to optimize and improve the faithfulness of the policy model $\pi_{\theta}$.
@@
$\begin{aligned} \mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{c,q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref})\right],\end{aligned}$
$\begin{aligned} \mathcal{L}_{i}&=\min\left(w_{i}A_{i},\text{clip}(w_{i},1-\epsilon,1+\epsilon)A_{i}\right),\end{aligned}$
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$, $\pi_{\theta_{\text{old}}}$ is the policy before the update, $\pi_{\text{ref}}$ is the reference policy (i.e., the initial model), $\epsilon$ and $\beta$ are hyperparameters controlling the update step and divergence regularization and $A_{i}$ is computed using the normalized reward within the group. We use our synthesized QA data as training data, which is easily verifiable, so that we can apply GRPO and train LLMs using rule-based rewards. Also, employing the rule-based GRPO removes the need for humans to label preference data for training the reward model.
@@ 2.3 Reward Design
Having a well-designed reward is key to the effectiveness of RL training (kimiteam2025kimik15scalingreinforcement). To use easily verifiable short-form QA data to improve the faithfulness of LLMs, the most intuitive reward would be the accuracy reward, which can check if the generated responses match the ground truth answers. However, in our early experiments, we found that relying solely on short-form QA data and accuracy rewards fails to enhance the faithfulness of long-form response generation, as the models may over-optimize short-form generation and learn a false pattern. The tuned models tend to copy text spans from the context as answers and lose their ability to generate long-form responses. Meanwhile, directly evaluating the faithfulness of long, free-form responses via the rule-based verification continues to pose an unresolved challenge.
Therefore, we propose Dual-GRPO, which includes a set of well-designed rewards that provide more harmonized guidance for optimizing LLMs to generate faithful responses. Unlike the original GRPO that over-optimizes short-form generation, we first prompt LLMs to generate both long-form and short-form responses, then assign different rewards to the two generated responses to improve the faithfulness.
System Prompt and Rollouts. For the provided context and question, we introduce a system prompt that requires LLMs to produce a reasoning process, then a long-form answer composed of detailed and complete sentences, and finally a concise short-form answer in just a few words. For example, given the context, if the question is “What is the country of origin of Super Mario?”, the long answer could be “Super Mario originated from Japan.”, while the short answer could simply be “Japan”. In this way, we can assign different reward scores to long-form and short-form answers while optimizing them both at once. This system prompt also triggers zero-shot chain-of-thought reasoning in the policy model, which progressively improves as training advances to optimize for the reward. The system prompt used for Dual-GRPO rollouts is shown in the Appendix B.
@@
Proxy Reward for Long-form Response Generation. Evaluating the faithfulness of long-form responses via the rule-based verification remains challenging. This is because these long-form answers are often free-form, making rule-based verification ineffective (openai2025deep). Therefore, instead of directly evaluating the faithfulness of the long-form response, we propose a proxy reward to evaluate it implicitly, as the faithfulness of a long-form answer can be measured by its ability to drive the LLMs toward a correct short-form answer. Specifically, for each generated long-form answer $y_{lf}$, we replace the given context $c$ with it as new input and feed it to the LLM to check whether the LLM can produce the correct short-form answer based on $y_{lf}$. If the generated long-form response can enable the LLM to generate the correct answer, it indicates that the long-form response stays faithful to the context, contains complete and easy-to-understand sentences, and correctly addresses the question. Thus, we assign a reward score of 1 for the positive long-form response that helps the LLM to produce the correct final answer, and a reward score of 0 for those that lead to an incorrect answer.
Format Reward. We also include a format reward that encourages adherence to a predefined output structure (e.g., using <think>, <long_answer>, and <short_answer>tags). Outputs that conform to this pattern receive a reward boost, thereby enhancing consistency. We use string matching to evaluate whether the generated responses adhere to the format, giving a score of 1 for a match and 0 for a mismatch.
Finally, we use the sum of these three rewards as the final composite reward. It enhances the efficacy of the RL training framework, guiding the model toward generating more faithful responses in both short-form and long-form tasks. More details are shown in the Appendix B.